Calculate the root-mean-square speed of molecular chlorine in m/s at 23.0 degrees Celsius.
Answer #1

Given temperature, 7= 23°C = (23 +213)K= 296K (TIK = TC +273) Molecule given: Molecular chlorine: cl Molar mass of = 71 g/mol